OpenStudy (anonymous):

Solve the given system, using the substitution method. y = 3x – 7 6x – 2y = 12 A. (14, 12) B. There are an infinite number of solutions. C. (12, 14) D. There is no solution.

jimthompson5910 (jim_thompson5910):

6x – 2y = 12 6x – 2(3x-7) = 12 ... replace y with 3x-7 (since y = 3x-7) 6x – 6x + 14 = 12 I'll let you finish up

OpenStudy (anonymous):

Thanks I just didnt get the 2(3x-7) part

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (abb0t):

Well, well, well...

OpenStudy (anonymous):

6x - 6x = 0. 12 = 14 what does that tell you?

OpenStudy (anonymous):

There is no solution. If you rearrange both of them in terms of y you'll see they both have the same slope and hence are parallel. Which means they don't intersect. try and plot them you'll see.

OpenStudy (anonymous):

oh ok thnx

OpenStudy (anonymous):

alot i was stumped